# Mode analysis of Nambu-Goldstone modes in U(1) charged first-order relativistic viscous hydrodynamics

### v2

@@ 2.1 Generating functional of UV model
These are the most general forms in terms of the symmetries associated with the external fields. All first-order coefficients are mass dimension three, which implies the effective theory is an expansion in $\varepsilon\equiv(\rho+P)^{-1}v\nabla$ with the viscosity coefficients $v=\epsilon,\lambda,\zeta$, etc…
@@
Eq. (8) is obtained by integrating the fast modes in the UV theory in the Schwinger-Keldysh formalism. In the operator formalism, the path integral should be written as
$e^{iI}={\rm Tr}[\hat{\varrho}\hat{U}^{\dagger}_{2}\hat{U}_{1}],$ (23)
@@
where $\hat{\varrho}$ is the density operator, and $\hat{U}_{i}$ is the unitary operator defined on the $i$th path in the closed-time-path. The external fields are suppressed for notational simplicity. The physical and noise fields in the Keldysh basis are defined as the average and difference of them defined on the first and second contours in the closed-time-path, such as $A\equiv(A_{1}+A_{2})/2$ and $\bar{A}\equiv A_{1}-A_{2}$. The consistency with the UV theory requires the following conditions Crossley:2015evo ; Glorioso:2017fpd for the generating functional:
  1. 1.
    Conjugate condition: the Hermitian condition for the density operator $\hat{\varrho}=\hat{\varrho}^{\dagger}$ leads to
    ${\rm Tr}[\hat{\varrho}\hat{U}^{\dagger}_{2}\hat{U}_{1}]^{\dagger}={\rm Tr}[\hat{U}^{\dagger}_{1}\hat{U}_{2}\hat{\varrho}^{\dagger}]={\rm Tr}[\hat{\varrho}\hat{U}^{\dagger}_{1}\hat{U}_{2}].$ (24)
@@
    $-I^{*}[g,A,\bar{g},\bar{A}]=I[g,A,-\bar{g},-\bar{A}].$ (25)
  2. 2.
    Unitarity condition: the Schwinger-Keldysh path integral is trivial if $\hat{U}_{1}=\hat{U}_{2}$. In the Keldysh basis, this corresponds to the absence of the noise:
    $I[g,A,\bar{g}=0,\bar{A}=0]=0.$ (26)
    The trace inequality $|{\rm Tr}[\varrho\hat{U}_{2}^{\dagger}\hat{U}_{1}]|\leq 1$ further implies (see also appendix of Ref. Ota:2024mps )
    ${\rm Im}~{}I\geq 0.$ (27)
  3. 3.
    Local KMS condition: Let $\hat{\mathcal{U}}_{i}=\mathcal{T}\hat{U}_{i}\mathcal{T}^{\dagger}$ be the time reversal transformation for the unitary operators, and the initial state satisfies $\rho=\mathcal{T}^{\dagger}\rho\mathcal{T}$. Then, the path integral is said to be symmetric under the micro-time reversal if
    ${\rm Tr}\left[\hat{\rho}\hat{U}_{1}\hat{U}_{2}^{-1}\right]={\rm Tr}\left[\hat{\rho}\hat{\mathcal{U}}_{2}^{-1}\hat{\mathcal{U}}_{1}\right].$ (28)
    The left hand side means that $\rho$ is the final condition for the path integral and the closed-time-path is reversed. The right hand side implies that arbitrary physical quantities in the path integral are time-reversed. For a thermal state, we can recast Eq. (28) into a more convenient form known as the local KMS condition, where the transformation is expressed by a $Z_{2}$ transformation of the external fields.
    First, let us write a thermal density operator
    $\hat{\varrho}=\frac{1}{Z}e^{-\beta(\hat{H}-\mu\hat{Q})},~{}Z\equiv{\rm Tr}[e^{-\beta(\hat{H}-\mu\hat{Q})}],$ (29)
    with the inverse temperature $\beta$, chemical potential $\mu$, Hamiltonian $\hat{H}$ and the conserved charge $\hat{Q}$. As $e^{-\beta\hat{H}}$ can be regarded as an imaginary time translation, we find
    ${\rm Tr}[e^{-\beta\hat{H}}U^{-1}_{2}U_{1}]$ $={\rm Tr}[e^{-\beta\hat{H}}U^{-1}_{2}e^{(\beta-d)\hat{H}}e^{-\beta\hat{H}}e^{d\hat{H}}U_{1}e^{-d\hat{H}}e^{d\hat{H}}]$ (30)
    where $d\in[0,\beta)$. Note that $\hat{Q}$ is suppressed since $[\hat{H},\hat{Q}]=0$. Using the trace cyclicity, we obtain
    ${\rm Tr}[e^{-\beta\hat{H}}U^{-1}_{2}U_{1}]={\rm Tr}[e^{-\beta\hat{H}}U_{1\uparrow}U^{-1}_{2\downarrow}],$ (31)
@@
    $U_{1\uparrow}$ $=e^{d\hat{H}}U_{1}e^{-d\hat{H}},$ (32)
    $U_{2\downarrow}^{-1}$ $=e^{-(\beta-d)\hat{H}}U_{2}^{-1}e^{(\beta-d)\hat{H}}.$ (33)
    For simplicity, we put $d=\beta/2$. Then, Eqs. (32) and (33) are equivalent to consider the following transformations for the external fields:
    $e^{\frac{i}{2}\hbar\pounds_{\beta}}A_{1}$ $=A_{1}+\frac{i\hbar}{2}\pounds_{\beta}A_{1}+\mathcal{O}(\hbar^{2}),$ (34)
    $e^{-\frac{i}{2}\hbar\pounds_{\beta}}A_{2}$ $=A_{2}-\frac{i\hbar}{2}\pounds_{\beta}A_{2}+\mathcal{O}(\hbar^{2}),$ (35)
    where £$\pounds_{\beta}$ is the Lie derivative along $\beta^{\mu}$, which generates the exponential map along the integral curve of $\beta^{\mu}$. Thus, a translation in the system reduces to the translation of the source. The finite non-diagonal imaginary time shift for $A_{1}$ and $A_{2}$ in the Keldysh basis is
    $A$ $\to A+\mathcal{O}(\hbar^{2}),$ (36)
    $\bar{A}$ $\to\bar{A}+i\hbar\pounds_{\beta}A+\mathcal{O}(\hbar^{3}).$ (37)
    The same discussion applies to $g_{i}$. We keep $\beta$ finite but consider the small $\hbar$ expansion. In this case, one can approximate the non-diagonal time translation by the noise time diffeomorphism generated by $i\hbar\beta^{\mu}$ in the Keldysh basis as above. The combination of transformations (28) and (31) reduces to a symmetry under a transformation for the external fields, which is known as local KMS symmetry. The mirco-time reversal symmetry $\mathcal{T}$ we imposed in the path integral can be generalized to $\mathcal{P}\mathcal{T}$, or $\mathcal{C}\mathcal{P}\mathcal{T}$ with parity flip ($\mathcal{P}$), and charge conjugation ($\mathcal{C}$).
Eq. (8) satisfies the conjugate condition (24) and the first unitarity condition (26) by construction. In the following section, we impose the second unitarity condition (27) and the local KMS condition for further constraining the undetermined coefficients in the constitutive relation.
@@ 2.2 Local KMS condition
In this subsection, we derive a consequence of the KMS condition. Consider a local KMS transformation: a $\mathcal{P}\mathcal{T}$ transformation followed by the noise time diffeomorphism with respect to $i\hbar\beta^{\mu}$. A covariant derivative and a coordinate system are $\mathcal{P}\mathcal{T}$-odd, and 4-velocity component $u^{\mu}$, the external fields $A_{\mu}$ and $g_{\mu\nu}$ are $\mathcal{P}\mathcal{T}$-even (see Appendix C of Ref. Liu:2018kfw ). Note that $u^{\mu}\partial_{\mu}$ is $\mathcal{P}\mathcal{T}$-odd and hence $u^{\mu}$ is even. Since $u^{\mu}$ is $\mathcal{P}\mathcal{T}$-even and $\nabla$ is $\mathcal{P}\mathcal{T}$-odd, $T_{(0)}^{\mu\nu}$ and $J_{(0)}^{\mu}$ are $\mathcal{P}\mathcal{T}$-even and $T_{(1)}^{\mu\nu}$ and $J_{(1)}^{\mu}$ are $\mathcal{P}\mathcal{T}$-odd. Thus, the $\mathcal{P}\mathcal{T}$ transformations for the non-fundamental quantities are not trivially given by the coordinate transformation, just like the parity transformation for the electric and magnetic fields.
The local KMS transformation of Eq. (8) is
$T^{\mu\nu}\bar{g}_{\mu\nu}$ $\to T_{(0)}^{\mu\nu}\bar{g}_{\mu\nu}-T_{(1)}^{\mu\nu}\bar{g}_{\mu\nu}+i\hbar T_{(0)}^{\mu\nu}\pounds_{\beta}g_{\mu\nu}-i\hbar T_{(1)}^{\mu\nu}\pounds_{\beta}g_{\mu\nu},$ (38)
$J^{\mu}\bar{A}_{\mu}$ $\to J^{\mu}_{(0)}\bar{A}_{\mu}-J^{\mu}_{(1)}\bar{A}_{\mu}+J^{\mu}_{(0)}i\hbar\pounds_{\beta}A_{\mu}-J^{\mu}_{(1)}i\hbar\pounds_{\beta}A_{\mu},$ (39)
$iW_{(0)}^{\mu\nu\alpha\beta}\bar{g}_{\mu\nu}\bar{g}_{\alpha\beta}$ $\to iW_{(0)}^{\mu\nu\alpha\beta}\left(\bar{g}_{\mu\nu}+i\hbar\pounds_{\beta}g_{\mu\nu}\right)\left(\bar{g}_{\alpha\beta}+i\hbar\pounds_{\beta}g_{\alpha\beta}\right),$ (40)
$iX^{\mu\nu\rho}_{(0)}\bar{g}_{\mu\nu}\bar{A}_{\rho}$ $\to iX^{\mu\nu\rho}_{(0)}\left(\bar{g}_{\mu\nu}+i\hbar\pounds_{\beta}g_{\mu\nu}\right)\left(\bar{A}_{\rho}+i\hbar\pounds_{\beta}A_{\rho}\right),$ (41)
$iY^{\mu\rho}_{(0)}\bar{A}_{\mu}\bar{A}_{\rho}$ $\to iY^{\mu\rho}_{(0)}(\bar{A}_{\mu}+i\hbar\pounds_{\beta}A_{\mu})(\bar{A}_{\rho}+i\hbar\pounds_{\beta}A_{\rho}).$ (42)
Note that we keep $\beta^{\mu}$ finite and consider the leading order contribution in the small $\hbar$ limit. The noise fields are counted as $\mathcal{O}(\hbar)$ since they are zero for the classical path. Then, the variation of the generating functional is
@@
$-\left(T_{(1)}^{\mu\nu}+\frac{\hbar}{2}W_{(0)}^{\mu\nu\alpha\beta}\pounds_{\beta}g_{\alpha\beta}+\hbar X^{\mu\nu\rho}_{(0)}\pounds_{\beta}A_{\rho}\right)\left(\bar{g}_{\mu\nu}+\frac{i}{2}\hbar\pounds_{\beta}g_{\mu\nu}\right)$ (43)
$-2\left(J^{\mu}_{(1)}+\hbar Y^{\mu\rho}_{(0)}\pounds_{\beta}A_{\rho}+\frac{\hbar}{2}X^{\rho\nu\mu}_{(0)}\pounds_{\beta}g_{\rho\nu}\right)\left(\bar{A}_{\mu}+\frac{i}{2}\hbar\pounds_{\beta}A_{\mu}\right)\bigg{]}.$ (44)
The first line is recast into the total derivative as follows. Firstly, we find
@@
$=(\rho+P)u^{\mu}u^{\nu}\nabla_{\mu}(\beta u_{\nu})+P\nabla_{\mu}\beta^{\mu}$
$=-(\rho+P)u^{\mu}\nabla_{\mu}\beta+P\nabla_{\mu}\beta^{\mu}$
$=-(\rho+P)u^{\mu}\nabla_{\mu}\beta-\beta^{\mu}\nabla_{\mu}P+\nabla_{\mu}(P\beta^{\mu})$
$=-\left[\beta\left(\frac{\partial P}{\partial\beta}\right)_{\mu}+\rho+P\right]u^{\mu}\nabla_{\mu}\beta-\beta\left(\frac{\partial P}{\partial\mu}\right)_{\beta}u^{\mu}\nabla_{\mu}\mu+\nabla_{\mu}(P\beta^{\mu})$ (45)
where we used
$dP(\beta,\mu)=\left(\frac{\partial P}{\partial\beta}\right)_{\mu}d\beta+\left(\frac{\partial P}{\partial\mu}\right)_{\beta}d\mu,$ (46)
and the fact that $\beta^{\mu}\nabla_{\mu}$ is a total derivative with respect to a parameter of the integral curve of $\beta^{\mu}$. In fact,
$\beta^{\mu}\nabla_{\mu}P$ $=\left(\frac{\partial P}{\partial\beta}\right)_{\mu}\beta^{\mu}\nabla_{\mu}\beta+\left(\frac{\partial P}{\partial\mu}\right)_{\beta}\beta^{\mu}\nabla_{\mu}\mu$
$=\beta\left(\frac{\partial P}{\partial\beta}\right)_{\mu}u^{\mu}\nabla_{\mu}\beta+\beta\left(\frac{\partial P}{\partial\mu}\right)_{\beta}u^{\mu}\nabla_{\mu}\mu.$ (47)
Also, one finds
$J^{\mu}_{(0)}\pounds_{\beta}A_{\mu}$ $=nu^{\mu}\beta^{\rho}\nabla_{\rho}A_{\mu}+nu^{\mu}A_{\rho}\nabla_{\mu}\beta^{\rho}$
$=n\beta^{\rho}(u^{\mu}\nabla_{\rho}A_{\mu})+nu^{\mu}A_{\rho}\nabla_{\mu}(\beta u^{\rho})$
$=n\beta^{\rho}\nabla_{\rho}(u^{\mu}A_{\mu})-n\beta^{\rho}(A_{\mu}\nabla_{\rho}u^{\mu})+nA_{\rho}u^{\rho}u^{\mu}\nabla_{\mu}\beta+nu^{\mu}A_{\rho}\beta\nabla_{\mu}u^{\rho}$
$=n\mu u^{\mu}\nabla_{\mu}\beta+n\beta^{\rho}\nabla_{\rho}\mu-n\beta^{\rho}A_{\mu}\nabla_{\rho}u^{\mu}+n\beta^{\mu}A_{\rho}\nabla_{\mu}u^{\rho}$
$=n\mu u^{\mu}\nabla_{\mu}\beta+n\beta^{\rho}\nabla_{\rho}\mu.$ (48)
To summarize, we obtain
$\frac{1}{2}T_{(0)}^{\mu\nu}\pounds_{\beta}g_{\mu\nu}+J^{\mu}_{(0)}\pounds_{\beta}A_{\mu}$
$=-\left[\beta\left(\frac{\partial P}{\partial\beta}\right)_{\mu}+\rho+P-\mu n\right]u^{\mu}\nabla_{\mu}\beta+\left[n-\left(\frac{\partial P}{\partial\mu}\right)_{\beta}\right]\beta^{\mu}\nabla_{\mu}\mu+\nabla_{\mu}(P\beta^{\mu}).$ (49)
With the thermodynamic relations (272), we obtain
$\frac{1}{2}T_{(0)}^{\mu\nu}\pounds_{\beta}g_{\mu\nu}+J^{\mu}_{(0)}\pounds_{\beta}A_{\mu}=\nabla_{\mu}(P\beta^{\mu}).$ (50)
@@
$T_{(1)}^{\mu\nu}+\hbar X^{\mu\nu\rho}_{(0)}\pounds_{\beta}A_{\rho}+\frac{\hbar}{2}W_{(0)}^{\mu\nu\alpha\beta}\pounds_{\beta}g_{\alpha\beta}$ $=0,$ (51)
$J^{\mu}_{(1)}+\hbar Y^{\mu\rho}_{(0)}\pounds_{\beta}A_{\rho}+\frac{\hbar}{2}X^{\rho\nu\mu}_{(0)}\pounds_{\beta}g_{\rho\nu}$ $=0,$ (52)
are satisfied, we find
$\delta_{\rm KMS}I=$ $\int d^{4}x\sqrt{-\det g}\nabla_{\mu}(P\beta^{\mu})+\mathcal{O}(\hbar^{3}),$ (53)
The boundary term is irrelevant to the variation with respect to the external sources, so it is absorbed in the normalization factor of the generating functional. When imposing the KMS condition, the noise terms reduce $\hbar$, which represents the classical statistical noise.

### v3

@@ 2.1 Generating functional of UV model
These are the most general forms allowed by general covariance. All first-order coefficients have mass dimension three. The effective theory is organized as an expansion in $\varepsilon\equiv(\rho+P)^{-1}v\nabla$, with viscosity coefficients $v=\epsilon,\lambda,\zeta$, etc. Readers can find a table summarizing the notational differences from a seminal work [23] in Appendix A.
2.2 Constraints on the generating functional
The above considerations align with the traditional derivative expansion of first-order hydrodynamics. With this setup, one can write the hydrodynamical equations and discuss the dynamics of $T^{\mu\nu}$ and $J^{\mu}$. Moreover, by constructing the entropy current, the local second law of thermodynamics introduces additional constraints on the hydrodynamical coefficients.
We reconstruct the effective action from these constitutive relations. The alternative to the local second law is the imposition of the unitarity and KMS conditions on the generating functional. With these conditions, the effective action is fixed up to the noise order, i.e., $W$, $X$, and $Y$ are related to $T$ and $J$.
@@ 2.2 Constraints on the generating functional
Eq. (2.8) is obtained by integrating out the fast modes in the UV theory using the Schwinger-Keldysh formalism. In the operator formalism, the path integral can be written as
$e^{iI}={\rm Tr}[\hat{D}\hat{U}^{\dagger}_{2}\hat{U}_{1}],$ (2.23)
@@
where $\hat{D}$ is the density operator, and $\hat{U}_{i}$ is the unitary operator defined on the $i$-th path of the closed time path (CTP). External fields are suppressed for notational simplicity. The physical and noise fields in the Keldysh basis are defined as the average and difference of those defined on the first and second contours of the CTP, namely $A\equiv(A_{1}+A_{2})/2$ and $\bar{A}\equiv A_{1}-A_{2}$. Consistency with the UV theory requires the generating functional to satisfy the following conditions [15, 24]:
Conjugate condition.
The Hermiticity of the density operator $\hat{D}=\hat{D}^{\dagger}$ implies
${\rm Tr}[\hat{D}\hat{U}^{\dagger}_{2}\hat{U}_{1}]^{\dagger}={\rm Tr}[\hat{U}^{\dagger}_{1}\hat{U}_{2}\hat{D}^{\dagger}]={\rm Tr}[\hat{D}\hat{U}^{\dagger}_{1}\hat{U}_{2}].$ (2.24)
@@
$-I^{*}[g,A,\bar{g},\bar{A}]=I[g,A,-\bar{g},-\bar{A}].$ (2.25)
Unitarity condition.
The Schwinger-Keldysh path integral yields a trivial result when $\hat{U}_{1}=\hat{U}_{2}$. In the Keldysh basis, this corresponds to the absence of noise:
$I[g,A,\bar{g}=0,\bar{A}=0]=0.$ (2.26)
The trace inequality $|{\rm Tr}[\hat{D}\hat{U}_{2}^{\dagger}\hat{U}_{1}]|\leq 1$ further implies (see also Appendix of Ref. [25])
${\rm Im}~I\geq 0.$ (2.27)
Local KMS condition.
The local KMS condition arises as a remnant of microscopic time-reversal symmetry in a thermal ensemble and plays a crucial role in constraining the structure of the effective action.
Let us first consider the condition for microscopic time-reversal symmetry. Define the time-reversed evolution operators as $\hat{\mathcal{U}}_{i}=\mathcal{T}\hat{U}_{i}\mathcal{T}^{\dagger}$, where $\mathcal{T}$ is a time-reversal operator. If the initial density matrix is invariant under time reversal, i.e., $\hat{D}=\mathcal{T}^{\dagger}\hat{D}\mathcal{T}$, then the following identity holds:
${\rm Tr}\left[\hat{D}\hat{U}_{1}\hat{U}_{2}^{-1}\right]={\rm Tr}\left[\hat{D}\hat{\mathcal{U}}_{2}^{-1}\hat{\mathcal{U}}_{1}\right].$ (2.28)
The left-hand side represents a CTP evolution ending in state $\hat{D}$, while the right-hand side corresponds to a time-reversed evolution of all operators. This microscopic symmetry implies a constraint on the generating functional, which underlies the so-called local KMS condition.
To understand its concrete implications, consider a system in thermal equilibrium with inverse temperature $\beta$ and chemical potential $\mu$. The corresponding density operator is
$\hat{D}=\frac{1}{Z}e^{-\beta(\hat{H}-\mu\hat{Q})},\quad Z={\rm Tr}[e^{-\beta(\hat{H}-\mu\hat{Q})}].$ (2.29)
For simplicity, we focus on the Hamiltonian part and assume $[\hat{H},\hat{Q}]=0$. Since $e^{-\beta\hat{H}}$ generates imaginary-time translations, the trace
${\rm Tr}[e^{-\beta\hat{H}}\hat{U}_{2}^{-1}\hat{U}_{1}]$ (2.30)
can be rewritten by inserting imaginary-time translation operators before and after $\hat{U}_{1}$ and $\hat{U}_{2}^{-1}$:
$={\rm Tr}[e^{-\beta\hat{H}}\hat{U}_{2}^{-1}e^{(\beta-d)\hat{H}}e^{-\beta\hat{H}}e^{d\hat{H}}\hat{U}_{1}e^{-d\hat{H}}e^{d\hat{H}}],$ (2.31)
where $d\in[0,\beta)$ is arbitrary. By cyclicity of the trace, this becomes
${\rm Tr}[e^{-\beta\hat{H}}\hat{U}_{1\uparrow}\hat{U}_{2\downarrow}^{-1}],$ (2.32)
@@
$\hat{U}_{1\uparrow}$ $=e^{d\hat{H}}\hat{U}_{1}e^{-d\hat{H}},$ (2.33)
$\hat{U}_{2\downarrow}^{-1}$ $=e^{-(\beta-d)\hat{H}}\hat{U}_{2}^{-1}e^{(\beta-d)\hat{H}}.$ (2.34)
Now, suppose the effective action is written in terms of external sources $A_{i}$ on the two legs of the CTP ($i=1,2$). The above imaginary-time translations correspond to a shift of these sources. For $d=\beta/2$, the infinitesimal form of this transformation becomes:
$e^{\pm\frac{i}{2}\hbar\pounds_{\beta}}A_{i}=A_{i}\pm\frac{i\hbar}{2}\pounds_{\beta}A_{i}+\mathcal{O}(\hbar^{2}),$ (2.35)
where £$\pounds_{\beta}$ is the Lie derivative along the thermal vector $\beta^{\mu}$. In the Keldysh basis,
$A$ $\to A+\mathcal{O}(\hbar^{2}),$ (2.36)
$\bar{A}$ $\to\bar{A}+i\hbar\pounds_{\beta}A+\mathcal{O}(\hbar^{3}),$ (2.37)
i.e., the physical field $A$ remains unchanged at leading order, while the noise field $\bar{A}$ acquires a shift proportional to £$\pounds_{\beta}A$.
This shift reflects a symmetry of the effective action under a combination of imaginary-time translations and time reversal, which is called the local KMS symmetry. It provides a $Z_{2}$-type constraint on the effective action that becomes manifest in the semiclassical ($\hbar\to 0$) limit. This symmetry is instrumental in deriving fluctuation-dissipation relations and ensuring the thermal consistency of dissipative dynamics.
While we have focused here on time-reversal symmetry $\mathcal{T}$, the same argument can be extended to $\mathcal{P}\mathcal{T}$ or $\mathcal{C}\mathcal{P}\mathcal{T}$ symmetry, depending on the discrete symmetries of the underlying UV theory.
Eq. (2.8) satisfies the conjugate condition (2.24) and the first unitarity condition (2.26) by construction. In the next section, we impose the second unitarity condition (2.27) and the local KMS condition to further constrain the undetermined coefficients in the constitutive relations.
@@ 2.3 Local KMS condition
In this subsection, we derive a consequence of the KMS condition. Consider a local KMS transformation, which consists of a $\mathcal{P}\mathcal{T}$ transformation followed by an infinitesimal noise time diffeomorphism generated by $i\hbar\beta^{\mu}$. A covariant derivative and the coordinate system are $\mathcal{P}\mathcal{T}$-odd, whereas the 4-velocity $u^{\mu}$ and the external fields $A_{\mu}$ and $g_{\mu\nu}$ are $\mathcal{P}\mathcal{T}$-even (see Appendix C of Ref. [26]). Note that $u^{\mu}\partial_{\mu}$ is $\mathcal{P}\mathcal{T}$-odd, implying that $u^{\mu}$ itself is even.
Since $u^{\mu}$ is $\mathcal{P}\mathcal{T}$-even and $\nabla$ is $\mathcal{P}\mathcal{T}$-odd, the zeroth-order quantities $T_{(0)}^{\mu\nu}$ and $J_{(0)}^{\mu}$ are $\mathcal{P}\mathcal{T}$-even, while the first-order corrections $T_{(1)}^{\mu\nu}$ and $J_{(1)}^{\mu}$ are $\mathcal{P}\mathcal{T}$-odd. Thus, the $\mathcal{P}\mathcal{T}$ transformations for these composite quantities are not trivially dictated by coordinate transformation, much like the parity transformation for the electric and magnetic fields.
The transformation of Eq. (2.8) under the local KMS symmetry is:
$T^{\mu\nu}\bar{g}_{\mu\nu}$ $\to T_{(0)}^{\mu\nu}\bar{g}_{\mu\nu}-T_{(1)}^{\mu\nu}\bar{g}_{\mu\nu}+i\hbar T_{(0)}^{\mu\nu}\pounds_{\beta}g_{\mu\nu}-i\hbar T_{(1)}^{\mu\nu}\pounds_{\beta}g_{\mu\nu},$ (2.38)
$J^{\mu}\bar{A}_{\mu}$ $\to J^{\mu}_{(0)}\bar{A}_{\mu}-J^{\mu}_{(1)}\bar{A}_{\mu}+J^{\mu}_{(0)}i\hbar\pounds_{\beta}A_{\mu}-J^{\mu}_{(1)}i\hbar\pounds_{\beta}A_{\mu},$ (2.39)
$iW_{(0)}^{\mu\nu\alpha\beta}\bar{g}_{\mu\nu}\bar{g}_{\alpha\beta}$ $\to iW_{(0)}^{\mu\nu\alpha\beta}\left(\bar{g}_{\mu\nu}+i\hbar\pounds_{\beta}g_{\mu\nu}\right)\left(\bar{g}_{\alpha\beta}+i\hbar\pounds_{\beta}g_{\alpha\beta}\right),$ (2.40)
$iX^{\mu\nu\rho}_{(0)}\bar{g}_{\mu\nu}\bar{A}_{\rho}$ $\to iX^{\mu\nu\rho}_{(0)}\left(\bar{g}_{\mu\nu}+i\hbar\pounds_{\beta}g_{\mu\nu}\right)\left(\bar{A}_{\rho}+i\hbar\pounds_{\beta}A_{\rho}\right),$ (2.41)
$iY^{\mu\rho}_{(0)}\bar{A}_{\mu}\bar{A}_{\rho}$ $\to iY^{\mu\rho}_{(0)}(\bar{A}_{\mu}+i\hbar\pounds_{\beta}A_{\mu})(\bar{A}_{\rho}+i\hbar\pounds_{\beta}A_{\rho}).$ (2.42)
We treat $\beta^{\mu}$ as finite but expand in small $\hbar$, keeping only the leading-order terms. Note that noise fields are counted as $\mathcal{O}(\hbar)$ since they vanish on the classical path.
The variation of the generating functional under this transformation is
@@
$-\left(T_{(1)}^{\mu\nu}+\frac{\hbar}{2}W_{(0)}^{\mu\nu\alpha\beta}\pounds_{\beta}g_{\alpha\beta}+\hbar X^{\mu\nu\rho}_{(0)}\pounds_{\beta}A_{\rho}\right)\left(\bar{g}_{\mu\nu}+\frac{i}{2}\hbar\pounds_{\beta}g_{\mu\nu}\right)$ (2.43)
$-2\left(J^{\mu}_{(1)}+\hbar Y^{\mu\rho}_{(0)}\pounds_{\beta}A_{\rho}+\frac{\hbar}{2}X^{\rho\nu\mu}_{(0)}\pounds_{\beta}g_{\rho\nu}\right)\left(\bar{A}_{\mu}+\frac{i}{2}\hbar\pounds_{\beta}A_{\mu}\right)\bigg{]}.$ (2.44)
The first line can be simplified to a total derivative. We first compute
@@
$=-\left[\beta\left(\frac{\partial P}{\partial\beta}\right)_{\mu}+\rho+P\right]u^{\mu}\nabla_{\mu}\beta-\beta\left(\frac{\partial P}{\partial\mu}\right)_{\beta}u^{\mu}\nabla_{\mu}\mu+\nabla_{\mu}(P\beta^{\mu})$ (2.45)
and
$J^{\mu}_{(0)}\pounds_{\beta}A_{\mu}=n\mu u^{\mu}\nabla_{\mu}\beta+n\beta^{\rho}\nabla_{\rho}\mu.$ (2.46)
Combining these results with the thermodynamical relations in App. B, we obtain
$\frac{1}{2}T_{(0)}^{\mu\nu}\pounds_{\beta}g_{\mu\nu}+J^{\mu}_{(0)}\pounds_{\beta}A_{\mu}=\nabla_{\mu}(P\beta^{\mu}).$ (2.47)
@@
$T_{(1)}^{\mu\nu}+\hbar X^{\mu\nu\rho}_{(0)}\pounds_{\beta}A_{\rho}+\frac{\hbar}{2}W_{(0)}^{\mu\nu\alpha\beta}\pounds_{\beta}g_{\alpha\beta}$ =0=0,= 0 , (2.48)
$J^{\mu}_{(1)}+\hbar Y^{\mu\rho}_{(0)}\pounds_{\beta}A_{\rho}+\frac{\hbar}{2}X^{\rho\nu\mu}_{(0)}\pounds_{\beta}g_{\rho\nu}$ =0=0,= 0 , (2.49)
are satisfied, the variation becomes
$\delta_{\rm KMS}I=$ $\int d^{4}x\sqrt{-\det g}\nabla_{\mu}(P\beta^{\mu})+\mathcal{O}(\hbar^{3}).$ (2.50)
The boundary term does not affect the variation of the generating functional with respect to external sources and can be absorbed into its normalization. Under the local KMS condition, noise terms consistently scale with $\hbar$, reflecting the emergence of classical statistical noise from quantum fluctuations. At this stage, we note that the local KMS condition can be generalized to a local vector $\beta^{\mu}(x)$. If the generating functional satisfies noise diffeomorphism invariance and micro-time reversality, one may introduce a local vector $\beta^{\mu}(x)$ that generates the thermodynamic relations and FDRs. This constitutes an implicit introduction of a local Gibbs distribution. The $Z_{2}$ transformation is expressed in a covariant manner.